OpenStudy (this):

hleppppp plllzz find the roots of the function x^2-10x+16=0 a,x=4,4 b,x=2,4 c,x=-2,8 d,x=-2,-8

hero (hero):

x^2 - 10x + 16 = 0 x^2 - 8x - 2x + 8*2 = 0 x(x-8) - 2(x-8) = 0 (x - 8)(x - 2) = 0 x - 8 = 0 x - 2 = 0 x = 8 x = 2

hero (hero):

Again, answer choices don't match

OpenStudy (this):

it wz b,x=2,8 i ve some kind of problem with copying down lol

OpenStudy (this):

thanks @Hero